1. At the beginning of the year, total assets amounted to P800,000 while total owner’s equity amounted to P550,000. During the year, assets increased by P50,000 while liabilities increased by P30%. If the owner contributed additional of P18,000 and withdrawn P5,000 during the year, how much is profit or loss earned or incurred by the entity? Indicate whether profit or loss.


1
Expert's answer
2021-10-05T13:33:49-0400

Total assets = Total liabilities + Total owner’s equity

P800,000 = Total liabilities + P550,000

Total liabilities = 800,000 - 550,000 = P250,000

Total assets (end of the year) = 800,000 + 50,000 = P850,000

Total liabilities (end of the year) = 250,000*(1+ 0.30) = P325,000

Total owner’s equity (end of the year) = 550,000 + 18,000 - 5,000 = P563,000

Total assets (end of year) = Total liabilities + Total owner’s equity

P850,000 = P325,000 + P563,000

P850,000 = P888,000

The amount of total liabilities and equity is greater than total assets that means loss incurred by the entity during the year.